OpenStudy (anonymous):

help please! i will give medal What is the value of the expression? A. 6 B. 9 C. 1 D. 5

OpenStudy (anonymous):

I'm guessing A

OpenStudy (anonymous):

@Asad786

OpenStudy (anonymous):

Are you in K12?

OpenStudy (anonymous):

yea, i just need help with this question, why

OpenStudy (anonymous):

I'm guessing its A

OpenStudy (anonymous):

I think its B but idk let me double check :) and Im in k12 too

OpenStudy (anonymous):

nice xD

OpenStudy (anonymous):

3 x 2 = 6 then add 6 + 6 | which is the top so it = 12

OpenStudy (anonymous):

then bottom:

OpenStudy (anonymous):

multiply comes first

OpenStudy (anonymous):

3 x 4 = 12 | then 12 - 10 = 2

OpenStudy (anonymous):

so its 12 / 2 = 6 | A

OpenStudy (anonymous):

you are right ! :)

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

thank you!